Grade 10Mechanics

The density of the material of cylindrical rod was determined by the formula d = m/pi r2L . The percentage error in m , r ,and l are 2% , 1.5% and 0.8% respectively . Calculate the maximum possible percentage error in determination of density.

Aiden Mathew Denny
8 Years agoGrade 10

3 Answers

Khimraj
ApprovedApproved Tutor Answer8 Years ago
density is given by
d = m/\pir2L
So
error maximum error is calculated as
\Deltad/d = \Deltam/m + 2\Deltar/r + \DeltaL/L 
error in density = 2 + 3 + 0.8 = 5.8%
Hope it clears. If you like answer then please approve it.
ajay kumar jha
6 Years ago
given : d=m/pir^2l and % error in m=2%,r=1.5% and l=0.8%
maximum error: deltad/d =delta m/m+2deltar/r+deltal/l= 2+1.5*2+0.8=5.8% 
hope it will be helpful for you
Vikas TU
6 Years ago
The density of the material of a cylindrical rod is determined by formula , D = m/πr²l
To calculate fractional error in D using formula,
∆D/D = ∆m/m + 2∆r/r + ∆l/l
and hence, % error in D = % error in m + 2 × % error in r + % error in l
given,
percentage error in m = 2%
percentage error in r = 1.5 %
percentage error in l = 0.8%
so, % error in D = 2 % + 2 × 1.5 % + 0.8%
= 2% + 3% + 0.8%
= 5.8%